OpenStudy (anonymous):

What is the equation of the line that passes through the point (6, 1) and is perpendicular to the line y = –3x + 2

OpenStudy (anonymous):

y=3x+2

OpenStudy (anonymous):

For perpendicular lines m1*m2=-1 y = –3x + 2 So m2=-1/-3=1/3 y=1/3 x+c when x=6, y=1 1=6/3+c 1=2+c c=-1 \[y=\frac{1}{3}x-1\]

OpenStudy (anonymous):

I have the options y=-3x+3 y=-3x-9 y=1/3x-1 y=1/3x+19

OpenStudy (anonymous):

When I put 1/3x-1 into my calculator they intersect??

OpenStudy (anonymous):

The answer is y=1/3 x-1. Yes they will intersect but the angle at which this happens is 90 degrees so it is perpendicular

OpenStudy (anonymous):

Ohhh okay :) Thank you so mcuh!!